# Variational Sparse Paired Autoencoders (vsPAIR) for Inverse Problems and Uncertainty Quantification

### v1

@@ Abstract
We propose the Variational Sparse Paired Autoencoder (vsPAIR) to address this challenge. The architecture pairs a standard VAE encoding observations with a sparse VAE encoding quantities of interest, connected through a learned latent mapping. The variational structure enables uncertainty estimation, the paired architecture encourages interpretability by anchoring QoI representations to clean data, and sparse encodings provide structure by concentrating information into identifiable factors rather than diffusing across all dimensions. We also propose modifications to existing sparse VAE methods: a hard-concrete spike-and-slab relaxation for differentiable training and a beta hyperprior for adaptive sparsity levels. To validate the effectiveness of our proposed architecture, we conduct experiments on blind inpainting and computed tomography, demonstrating that vsPAIR is a capable inverse problem solver that can provide interpretable and structured uncertainty estimates.
@@ 1 Introduction
More recently, there has been a shift towards data-driven approaches. Such approaches use parameterized neural networks to learn various attributes of the underlying inverse problem. End-to-end methods learn direct mappings $F^{\leftarrow}:\mathcal{Y}\rightarrow\mathcal{X}$ from paired samples $(x,y)$ [28, 59]. Prior learning methods employ a data-driven prior: plug-and-play methods incorporate pretrained denoisers as implicit priors within optimization schemes [55, 48], and learned methods use regularization techniques to replace hand-crafted penalties with trainable functionals [20, 40]. Hybrid methods such as algorithm unrolling interpret iterative solvers as neural networks with learnable parameters [3, 44]. Various other methodologies exist and have been successful in a variety of applications.
@@
The introduction of sparsity in the QoI autoencoder causes technical challenges. While our approach primarily utilizes the sVAE introduced in [54], we propose a series of modifications to this framework. In particular, we introduce a Bayesian variation of this approach which leverages a beta hyperprior to learn sparsity levels in the encoding process, as well as a hard-concrete spike-and-slab relaxation that builds off of existing approaches to enable exact-zeros in the latent representation with differentiable training.
@@
Section 2 reviews classical and learning-based formulations of inverse problems, including autoencoders and paired latent mappings. Section 3 introduces the vsPAIR framework: sVAE variations, the paired architecture, and theoretical analysis of uncertainty quantification under simplifying assumptions. Finally, Section 4 and Section 5 present numerical demonstrations on blind inpainting and computed tomography, followed by concluding remarks.
@@ 2 Background
This section reviews the classical formulations of inverse problems and the learning-based methods that motivate our approach.
@@ 2.1 Autoencoders and Variational Autoencoders (VAEs)
where $p(z)$ is an assumed prior over the latent space, typically $\mathcal{N}(z\mid 0,I)$ and ${\rm D}_{\rm KL}(\,\cdot\;||\;\cdot\,)$ denotes the KL-divergence [45]. Under Gaussian assumptions on the probabilistic encoder and decoder as well as unit Gaussian assumption on the prior, the reconstruction term reduces to scaled mean-squared error and the KL term admits a closed form; see [32] for a comprehensive tutorial. For further simplicity, we also assume that the encoder and decoder have diagonal and unit covariances respectively. Negating Equation 3, and plugging in these expressions, we obtain the VAE objective to be minimized in training:
@@ 2.2 Variational Sparse Autoencoders (sVAE)
Variational Sparse Autoencoders (sVAEs) extend the classical VAE framework by explicitly encouraging sparse latent representations. In a typical VAE, we use a standard Gaussian prior and learn to trade off reconstruction quality against a penalty that keeps the approximate posterior close to that prior. While this Gaussian-KL penalty does encourage structured latent representations, it will utilize all of the latent dimensions regardless of whether they hold relevant information or not. In contrast, a sVAE adopts an overcomplete latent representation, often substantially larger than that of a standard VAE, and explicitly promotes sparsity by encouraging many latent coordinates to be zero.
@@
Training of the sVAE must be handled carefully since the immediate question of how to choose a prior sparsity level $\rho$ arises and the Bernoulli sampling procedure further introduces complexity due to non-differentiability. These challenges are discussed in Section 3.1.
The benefit of sVAEs are that they typically lead to more interpretable, disentangled representations, smaller effective model capacity, and often better generalization on downstream tasks [54]. Because many real‐world signals lie on low‐dimensional manifolds, enforcing sparsity helps the model focus its capacity on the most meaningful latent factors, rather than spreading small amounts of information over all dimensions. They bridge the gap between classical sparse coding, which enforces strict zeros but lacks a probabilistic decoder, and VAEs, which offer a principled probabilistic formulation but no mechanism for efficient sparse representation.
@@ 2.3 Paired Autoencoder Frameworks (Latent Twins)
Introduced in [17], the latent twin methodology learns mathematical operators between latent representations of quantities of interest. This approach is broadly applicable to many areas of research including PDEs, ODEs, and inverse problems and can be adjusted to encode physical constraints on the system of interest in the learned latent space. To be precise, our work builds off of paired autoencoders for inverse problems, a precursor to the latent twin architecture.
@@ 3 Methodology: vsPAIR Framework
In this section, we present the key contribution of this work: the Variational Sparse Paired Autoencoder (vsPAIR). We begin by introducing two innovations that draw from previous work on model optimization with sparsity constraints and form a core component of vsPAIR enabling its uncertainty-aware paired architecture.
3.1 sVAE Innovations
Beta Hyperprior for Adaptive Sparsity.
Introducing sparsity into sVAE is not trivial and a sparsity level, $(1-\rho)$ for the prior must be specified. To address the latter issue, [54] originally leveraged a classifier-based approach. In order to take advantage of the whole latent-space for each embedding, we instead analyze this problem in a Bayesian sense, and assume that the distribution of $\rho$ is described by some underlying hyperprior distribution. In particular, we treat $\rho$ as a learnable parameter regularized by a Beta hyperprior
$\rho\sim\mathrm{Beta}(\alpha_{0},\beta_{0}).$ (15)
and consider the negative log-density as a training objective dependent on $\rho$,
$\mathcal{L}_{\rho}(\rho\,;\,\alpha_{0},\beta_{0})=\lambda_{\rho}\bigl[(\alpha_{0}-1)\log\rho+(\beta_{0}-1)\log(1-\rho)\bigr].$ (16)
Here, $\lambda_{\rho}$ controls the regularization strength, and terms constant in $\rho$ are dropped. The likelihood term seeks to push $\rho$ towards a probable value under the prescribed hyperprior, but is competing with the sVAE loss term. This establishes a tradeoff between sVAE performance and sparsity, for which optimization finds a balance.
Since $\rho$ represents the prior probability of a coordinate being active, larger $\rho$ corresponds more dense reconstructions. Setting $\alpha_{0}>\beta_{0}$ yields a left-skewed beta density that concentrates mass toward higher values of $\rho$, encouraging dense latent representations. Conversely, sparsity is favored when $\alpha_{0}<\beta_{0}$. During training, $\rho$ adapts to balance sparsity against reconstruction quality: for $\alpha_{0}<\beta_{0}$, if deactivating a dimension negligibly affects reconstruction, $\rho$ decreases to encourage more zeros; if reconstruction suffers, $\rho$ increases to preserve active dimensions. This self-tuning mechanism replaces manual hyperparameter selection with a principled Bayesian approach. Numerical results confirm that this approach aligns with these expectations–see Appendix A.
Hard-Concrete Relaxation.
A central challenge in training sparse VAEs arises from the non-differentiable Bernoulli sampling procedure: the loss involves quantities computed from draws of Bernoulli trials. [54] address this issue via a soft relaxation based on a scaled sigmoid. In particular, the latent encoding is computed element-wise as
$z_{i}=T(\eta_{i}-1+\omega_{i})\;\cdot\;(\mu_{i}+\sigma_{i}\epsilon)$ (17)
To enforce true sparsity, we adopt a hard-concrete relaxation that yields exact zeros in the forward pass while remaining compatible with gradient-based training. Our work applies several existing techniques to sVAE training–we use the smooth approximation for the Bernoulli random variable from [54, 41] and the straight-through derivative estimator from [7]. The gating scheme aligns with [39]. Here, we use soft-gate to refer to operations for which gradients flow through, and hard-gate for operations where the gradient is not tracked. To be precise, for each latent coordinate $i$, we proceed as follows:
  1. 1.
    Sample uniform noise $\eta_{i}\sim\mathrm{Uniform}(0,1)$.
  2. 2.
    Compute a soft-gate via the differentiable step function approximation: $s_{i}=\sigma\bigl(c(\eta_{i}-1+\omega_{i})\bigr)\in(0,1)$, where $c>0$ is a temperature parameter ([54, 41]).
  3. 3.
    Threshold to obtain a hard-gate: $\hat{s}_{i}=\mathbf{1}_{\{s_{i}>0.5\}}(s_{i})\in\{0,1\}$.
  4. 4.
    Construct a the final gate using a straight-through estimator: $\mathrm{gate}_{i}=\hat{s}_{i}-\mathrm{sg}(s_{i})+s_{i},$ where $\mathrm{sg}(\cdot)$ denotes stop-gradient. [7]
  5. 5.
    Apply the standard reparameterization ${z}_{i}^{*}=\mu_{i}+\sigma_{i}\epsilon_{i}$ with $\epsilon_{i}\sim\mathcal{N}(0,1)$, then mask: $z_{i}=\mathrm{gate}_{i}\cdot z_{i}^{*}$ [31].
At step 4, this gate takes binary values (from $\hat{s}_{i}$) but passes gradients through the soft approximation $s_{i}$. For backpropagation, neither the hard-gate that generates $\hat{s_{i}}$ nor the negative $s_{i}$ gradient is tracked. Thus, this approach produces hard thresholded values but only tracks gradients for the differentiable approximation of the step function: gradients flow through the soft-gate, but the hard-gate enforces exact zeros.
@@
Training follows the structure of the deterministic PAIR framework [23], with reconstruction losses replaced by the corresponding ELBOs and an additional term for the beta hyperprior. The complete objective is
$\mathcal{L}(\phi^{e_{x}},\theta^{d_{x}},\phi^{e_{y}},\theta^{d_{y}},\theta_{M},\rho)=\mathbb{E}_{(x,y)\sim p(x,y)}\big[\lambda_{1}\mathcal{L}_{\text{sVAE}}(\phi^{e_{x}},\theta^{d_{x}})+\lambda_{2}\mathcal{L}_{\text{VAE}}(\phi^{e_{y}},\theta^{d_{y}})+\lambda_{3}\mathcal{L}_{M}(\theta_{M})+\lambda_{\rho}\mathcal{L}_{\rho}(\rho\,;\,\alpha_{0},\beta_{0})\big],$ (18)
where $\mathcal{L}_{\text{sVAE}}$ (Equation 11) and $\mathcal{L}_{\text{VAE}}$ (Equation 4) are the ELBO losses for the QoI and observation autoencoders respectively, $\mathcal{L}_{M}$ is the latent mapping loss defined below, and $\mathcal{L}(\rho\,;\,\alpha_{0},\beta_{0})$ is the beta hyperprior term from (16). The non-negative hyperparameters $\lambda_{1},\ldots,\lambda_{\rho}$ control the relative weighting of each term.
We now address options for training the latent mapping. The first design choice that must be made is whether the autoencoders are to be trained first, and then the latent map after, or if should training proceed all together. If trained independently, then the latent codes for $X$ do not depend on $Y$ and a conditional independence assumption is embedded into the architecture: $Z_{x}$ will be conditionally independent on $Y$ given $X$; $Z_{y}$ will be conditionally independent on $X$ given $Y$. Of course, this conditional independence is imposed by the model architecture and need not hold for the true data distribution.
@@
The second design choice concerns the implementation of the latent mapping. For simplicity, we once again omit dependence on trainable parameters–that is, $\mu_{\phi_{x}}(x)$ is written as $\mu_{x}(x)$. We propose learning $M^{\leftarrow}$ as a direct mapping from the distribution parameters of the observation encoder to the distribution parameters of the QoI encoder: $M^{\leftarrow}$ learns to map $(\mu_{y}(y),\sigma_{y}(y))$ to $(\mu_{x}(x),\sigma_{x}(x),\omega_{x}(x))$ for $(x,y)\sim p(x,y)$. Here, inversion proceeds by first encoding the observation $y$, then computing approximate conditional distribution parameters for $x$, $(\hat{\mu}(y),\hat{\sigma}(y),\hat{\omega}(y))$, via $M^{\leftarrow}$, then decoding a sample $\hat{z}_{x}$ from this approximate distribution to obtain $\hat{x}$. Generation can be accomplished by sampling the approximate latent distribution and decoding. This method gives us an approximation of the posterior distribution $p(x\mid y)$ induced by the encoder conditionals, and we theoretically analyze a simplification of this approach in Section 3.3. $M^{\leftarrow}$ can be trained via MSE loss between the true and approximate parameters:
@@
There are many other ways to train this mapping, and many techniques from generative modeling are applicable. For example, one approach is to directly match the encoded latent samples by minimizing a reconstruction loss between $z_{x}$ and $M^{\leftarrow}(z_{y})$. If an MSE loss is used here, then $M^{\leftarrow}$ is such that
@@
In the case that $\ell_{x}=\ell_{y}$ have the same dimension, we may approach the training of $M^{\leftarrow}$ as a normalizing flow between samples drawn from $q_{\phi_{y}}(z_{y}\mid y)$ and $q_{\phi_{x}}(z_{x}\mid x)$ [53]. In this case, the pushforward of $q_{\phi_{y}}(z_{y}\mid y)$ under $M^{\to}$ is learned to approximate $p(z_{x}\mid x)$.
@@ 3.3 Uncertainty Quantification
Spike-and-slab distributions pose analytical challenges because the point mass at zero does not admit a Lebesgue density. For tractability, we instead reduce to the case where the variational posteriors for both $x$ and $y$ are Gaussian. That is, we assume that the latent representation for $x$, $z_{x}\sim q_{\phi_{x}}(z_{x}\mid x)=\mathcal{N}(z_{x}\mid\mu_{x}(x),\Sigma_{x}(x))$ and $z_{y}\sim q_{\phi_{y}}(z_{y}\mid y)=\mathcal{N}(z_{y}\mid\mu_{y}(y),\Sigma_{y}(y))$ where the covariances are symmetric positive definite for all $x$ and $y$.
@@
    $(X,Y)$ emits a joint density, $p(x,y)$
@@
  4. 4.
    The mapping from $x$ to the latent mean, $\mu_{x}(x)$ is affine, with $\mu_{x}(x)=Bx+c$, and $q_{\phi_{x}}(z_{x}\mid x)$, has constant symmetric positive definite covariance, $\Sigma_{x}(x)=\Sigma_{Z_{x}}$.
The first and second conditions are standard assumptions on the true data distribution, while the third can be induced by the model architecture: this is satisfied if the two encoders are trained independently. The fourth assumption is the most restrictive but enable an explicit formulation of the conditional network parameters in terms of the observation. Our main result is stated in Theorem 1 which establishes that the conditional mean and covariance of $Z_{x}\mid Y$ exist as a continuous function of $y$.
@@
Suppose that Assumptions 1, 2, 3 and 4 are satisfied. Then, the conditional mean and covariance of $Z_{x}\mid Y$ induced by the paired framework exist as continuous functions of $y$. In particular,
$\mathbb{E}_{Z_{x}\mid Y}[Z_{x}\mid Y]=B\mathbb{E}[X\mid Y]+c$ (21)
@@
The proof of this theorem follows from the conditional independence and towering property of conditional expectations, see Section C.1. The utility of Theorem 1 is that it tells us that our training objective for the latent mapping is well-posed in this simplistic setup: the mapping from encoded observation parameters $(\mu_{y}(y),\Sigma_{y}(y))$ to QoI posterior parameters is continuous as a function of $y$, hence learnable by the Universal Approximation Theorem assuming sufficient network expressiveness. This justifies training the latent mapping on paired samples $(\mu_{y}(y),\Sigma_{y}(y))$ and $(\mu_{x}(x),\Sigma_{x}(x))$ for $(x,y)\sim p(x,y)$.
@@
The proof, given in Section C.2, follows from marginalization and conditional independence. The utility of Theorem 2 is that under these conditions, the mapped parameters fully characterize $q(z_{x}\mid y)$: sampling an approximation $\hat{z}_{x}\sim\mathcal{N}(\hat{z}_{x}\mid\hat{\mu}(y),\hat{\Sigma}(y))$ with mapped parameters produces samples from the correct conditional distribution, not merely a distribution with matching moments.
Taken together, these results provide a theoretical foundation for the vsPAIR approach to uncertainty quantification under simplifying assumptions. Theorem 1 establishes the learnability of the latent mapping, while Theorem 2 guarantees correctness of the induced distribution in the linear–Gaussian setting. While these assumptions are restrictive, they isolate the essential mechanisms of the method and demonstrate that the proposed architecture is sound in principle. Its performance beyond this regime, including fully nonlinear problems, is investigated numerically in Section 4.

### v2

@@ Abstract
We propose the Variational Sparse Paired Autoencoder (vsPAIR) to address this challenge. The architecture pairs a standard VAE encoding observations with a sparse VAE encoding quantities of interest, connected through a learned latent mapping. The variational structure enables uncertainty estimation, the paired architecture encourages interpretability by anchoring QoI representations to clean data, and sparse encodings provide structure by concentrating information into identifiable factors rather than diffusing across all dimensions. To validate the effectiveness of our proposed architecture, we conduct experiments on blind inpainting and computed tomography, demonstrating that vsPAIR is a capable inverse problem solver that can provide interpretable and structured uncertainty estimates.
Keywords: Machine Learning, Inverse Problems, Uncertainty Quantification, Generative Modeling
@@ 1 Introduction
More recently, there has been a shift towards data-driven approaches. Such approaches use parameterized neural networks to learn various attributes of the underlying inverse problem. End-to-end methods learn direct mappings $F^{\leftarrow}:\mathcal{Y}\rightarrow\mathcal{X}$ from paired samples $(x,y)$ [31, 69]. Prior learning methods employ a data-driven prior: plug-and-play methods incorporate pretrained denoisers as implicit priors within optimization schemes [64, 55], and learned methods use regularization techniques to replace hand-crafted penalties with trainable functionals [22, 44]. Hybrid methods such as algorithm unrolling interpret iterative solvers as neural networks with learnable parameters [2, 49]. Various other methodologies exist and have been successful in a variety of applications [16].
@@
Section 2 reviews classical and learning-based formulations of inverse problems, including autoencoders and paired latent mappings. Section 3 introduces the vsPAIR framework: the paired architecture, and theoretical analysis of uncertainty quantification under simplifying assumptions. Finally, Section 4 and Section 5 present numerical demonstrations on blind inpainting and CT, followed by concluding remarks. The introduction of sparsity in the QoI autoencoder presents technical challenges which are discussed and addressed in Appendix A.
@@ 2 Background
This section reviews the classical formulations of inverse problems and the learning-based methods that motivate our approach. The reader is referred to [16] for a more in-depth exposition; we closely follow the structure and notation of this work.
@@ 2.1 Autoencoders and Variational Autoencoders (VAEs)
where $p(z)$ is an assumed prior over the latent space, typically $\mathcal{N}(z\mid 0,I)$, and ${\rm D}_{\rm KL}(\,\cdot\;||\;\cdot\,)$ denotes the KL-divergence [50]. Under Gaussian assumptions on the probabilistic encoder and decoder as well as unit Gaussian assumption on the prior, the reconstruction term reduces to scaled mean-squared error and the KL term admits a closed form; see [36] for a comprehensive tutorial. For further simplicity, we also assume that the encoder and decoder have diagonal and unit covariances respectively. Negating Equation 3, and plugging in these expressions, we obtain the VAE objective to be minimized in training:
@@ 2.2 Variational Sparse Autoencoders (sVAE)
Variational Sparse Autoencoders (sVAEs) extend the classical VAE framework by explicitly encouraging sparse latent representations. In a typical VAE, we use a unit Gaussian prior and learn to trade off reconstruction quality against a penalty that keeps the approximate posterior close to that prior. While this Gaussian-KL penalty does encourage structured latent representations, it will utilize all of the latent dimensions regardless of whether they hold relevant information or not. In contrast, a sVAE adopts an overcomplete latent representation, often substantially larger than that of a standard VAE, and explicitly promotes sparsity by encouraging many latent coordinates to be zero.
@@
Training of the sVAE must be handled carefully since the immediate question of how to choose a prior sparsity level $\rho$ arises and the Bernoulli sampling procedure further introduces complexity due to non-differentiability. These challenges are discussed in Appendix A, where we describe a training objective we use to learn $\rho$.
The benefit of sVAEs are that they typically lead to more interpretable, disentangled latent representations [62]. Because many real‐world signals lie on low‐dimensional manifolds, enforcing sparsity helps the model focus its capacity on the most meaningful latent factors, rather than spreading small amounts of information over all dimensions. They bridge the gap between classical sparse coding, which enforces strict zeros, and VAEs, which offer a probabilistic formulation but no mechanism for efficient sparse representation.
@@ 2.3 Paired Autoencoder Frameworks (Latent Twins)
Introduced in [18], the latent twin methodology learns mathematical operators between latent representations of quantities of interest. This approach is broadly applicable to many areas of research including PDEs, ODEs, and inverse problems and can be adjusted to encode physical constraints on the system of interest in the learned latent space. To be precise, our work builds off of paired autoencoders for inverse problems, a precursor to the latent twin architecture. Once again, we follow the notation and formulation used in [16].
@@ 3 Methodology: vsPAIR Framework
In this section, we present the key contribution of this work: the Variational Sparse Paired Autoencoder (vsPAIR). For details on the sVAE implementation, see Appendix A.
@@
Training follows the structure of the deterministic PAIR framework [25], with reconstruction losses replaced by the corresponding ELBOs. The complete objective is
$\mathcal{L}(\phi_{x},\theta_{x},\phi_{y},\theta_{y},\theta_{M},\rho)=\mathbb{E}_{(x,y)\sim p(x,y)}\big[\lambda_{1}\mathcal{L}_{\text{sVAE}}(\phi_{x},\theta_{x})+\lambda_{2}\mathcal{L}_{\text{VAE}}(\phi_{y},\theta_{y})+\lambda_{3}\mathcal{L}_{M}(\theta_{M})+\lambda_{\rho}\mathcal{L}_{\rho}(\rho\,;\,\alpha_{0},\beta_{0})\big],$ (15)
where $\mathcal{L}_{\text{sVAE}}$ (Equation 11) and $\mathcal{L}_{\text{VAE}}$ (Equation 4) are the ELBO losses for the QoI and observation autoencoders respectively, $\mathcal{L}_{M}$ is the latent mapping loss defined below and $\mathcal{L}(\rho\,;\,\alpha_{0},\beta_{0})$ is a beta hyperprior regularization used to learn the prior sparsity level–see Appendix A, Equation 21. The non-negative hyperparameters $\lambda_{1},\ldots,\lambda_{\rho}$ control the relative weighting of each term.
We now address options for training the latent mapping. The first design choice that must be made is whether the autoencoders are to be trained first, and then the latent map after, or if should training proceed all together. If trained independently, then the latent codes for $X$ do not depend on $Y$ and a conditional independence assumption is embedded into the architecture: $Z_{x}$ will be conditionally independent on $Y$ given $X$; $Z_{y}$ will be conditionally independent on $X$ given $Y$. Of course, this conditional independence is imposed by the model architecture and need not hold for the true underlying distribution.
@@
The second design choice concerns the implementation of the latent mapping. For simplicity, we once again omit dependence on trainable parameters–that is, $\mu_{\phi_{x}}(x)$ is written as $\mu_{x}(x)$. We propose learning $M^{\leftarrow}$ as a direct mapping from the distribution parameters of the observation encoder to the distribution parameters of the QoI encoder: $M^{\leftarrow}$ learns to map $(\mu_{y}(y),\sigma_{y}(y))$ to $(\mu_{x}(x),\sigma_{x}(x),\omega_{x}(x))$ for $(x,y)\sim p(x,y)$. Here, inversion proceeds by first encoding the observation $y$, then computing approximate conditional distribution parameters for $x$, $(\hat{\mu}(y),\hat{\sigma}(y),\hat{\omega}(y))$, via $M^{\leftarrow}$, then decoding a sample $\hat{z}_{x}$ from this approximate distribution to obtain $\hat{x}$. Generation can be accomplished by sampling the approximate latent distribution and decoding. This method provides an approximation of the posterior distribution $q(z_{x}\mid y)$ induced by the encoder conditionals, and we theoretically analyze a simplification of this approach in Section 3.2. $M^{\leftarrow}$ can be trained via MSE loss between the true and approximate parameters:
@@
There are many other ways to train this mapping, and many techniques from generative modeling are applicable. For example, one approach is to directly match the encoded latent samples by minimizing a reconstruction loss between $z_{x}$ and $M^{\leftarrow}(z_{y})$. If an MSE loss is used here, then $M^{\leftarrow}$ can approximate
@@
In the case that $\ell_{x}=\ell_{y}$ have the same dimension, we may approach the training of $M^{\leftarrow}$ as a normalizing flow between samples drawn from $q_{\phi_{y}}(z_{y}\mid y)$ and $q_{\phi_{x}}(z_{x}\mid x)$ [61]. In this case, the pushforward of $q_{\phi_{y}}(z_{y}\mid y)$ under $M^{\to}$ is learned to approximate $q_{\phi_{x}}(z_{x}\mid x)$.
@@ 3.2 Uncertainty Quantification
Spike-and-slab distributions pose analytical challenges because of the point mass at zero does not admit a Lebesgue density. For tractability, we instead reduce to the case where the variational posteriors for both $x$ and $y$ are Gaussian. That is, we assume that the latent representation for $x$, $z_{x}\sim q_{\phi_{x}}(z_{x}\mid x)=\mathcal{N}(z_{x}\mid\mu_{x}(x),\Sigma_{x}(x))$ and $z_{y}\sim q_{\phi_{y}}(z_{y}\mid y)=\mathcal{N}(z_{y}\mid\mu_{y}(y),\Sigma_{y}(y))$ where the covariances are symmetric positive definite for all $x$ and $y$.
@@
    $\mathbb{E}[X\mid Y=y]$ and ${\rm Cov}(X\mid Y=y)$ are both continuous as functions of $y$.
@@
    The mapping from $x$ to the latent mean is affine, $\mu_{x}(x)=Bx+c$, and the latent covariance mapping is constant, $\Sigma_{x}(x)=\Sigma_{Z_{x}}$.
The first condition is an assumption on the true data distribution, while the second can be induced by the model architecture: this is satisfied if the two encoders are trained independently. The third depends on the networks selected for the variational posterior for the QoI and is the most restrictive assumption. Our main result is stated in Theorem 1 which establishes that the conditional mean and covariance of $Z_{x}\mid Y$ exist as a continuous function of $y$.
@@
Suppose that Assumptions 1, 2, and 3 are satisfied. Then, the conditional mean and covariance of $Z_{x}\mid Y$ induced by the paired framework exist as continuous functions of $y$. In particular,
$\mathbb{E}[Z_{x}\mid Y]=B\mathbb{E}[X\mid Y=y]+c$ (18)
@@
The expressions follow directly from the towering property of conditional expectations and the law of total covariance [33], see Section C.1. The utility of Theorem 1 is that it tells us that our training objective for latent mapping is well-posed in this simplistic setup: the mapping from encoded observation parameters $(\mu_{y}(y),\Sigma_{y}(y))$ to QoI posterior parameters is continuous as a function of $y$, hence it can be learned by the Universal Approximation Theorem assuming sufficient network expressiveness and sufficient regularity on the expressions above [27]. This justifies training the latent mapping on paired samples of $(\mu_{y}(y),\Sigma_{y}(y))$ and $(\mu_{x}(x),\Sigma_{x}(x))$ for $(x,y)\sim p(x,y)$.
@@
The proof, given in Section C.2, follows from marginalization, conditional independence and results from Bayesian statistics. The utility of Theorem 2 is that under these conditions, the mapped parameters fully characterize $q(z_{x}\mid y)$: sampling an approximation $\hat{z}_{x}\sim\mathcal{N}(\hat{z}_{x}\mid\hat{\mu}(y),\hat{\Sigma}(y))$ with mapped parameters produces samples from the correct conditional distribution, not merely a distribution with matching moments.
Taken together, these results provide a theoretical foundation for the vsPAIR approach to uncertainty quantification under simplifying assumptions. Theorem 1 establishes that the desired latent mapping exists as an explicit function of the observations, while Theorem 2 guarantees correctness of the induced distribution in the linear–Gaussian setting. While these assumptions are restrictive, they isolate the essential mechanisms of the method and demonstrate that the proposed architecture is sound in principle. Its performance beyond this regime is investigated numerically in Section 4.